A man built a house costing 9 times as much as the lot on which it was built. If the lot cost $5,000 less than 1/6 the house. What was the cost of each?
Thanks very much for your help!

Answer by checkley77(12844)

H=9L
L=H/6-5,000
H=9(H/6-5,000)
H=9H/6-45,000
9H/6-H=45,000
H/2=45,000
H=90,000 COST OF THE HOUSE.
L=90,000/6-5,000
L=15,000-5,000
L=$10,000 COST OF THE LOT.
PROOF:
90,000=910,000
90,000=90,000